OpenStudy (anonymous):

A student opens a mathematics book to two facing pages. The product numbers of the page number is 156. Find the page numbers... this question confuses me??

OpenStudy (anonymous):

So there are two number let the first equal x and the second equal y ok?

OpenStudy (vishal_kothari):

x is left hand page x+1 is right hand page x(x+1)=156 x^2 +x=156 x^2+x-156=0 (x+13)(x-12 )=0 There are two possible solutions to the equation. x+13 = 0 and x-12=0 or x=-13 and x=12 since you can't have a negative page number x=12 for the lh page and x+1 or 13 for the right hand page..

OpenStudy (anonymous):

ok

OpenStudy (vishal_kothari):

kk

OpenStudy (anonymous):

Dont give answers bro

OpenStudy (vishal_kothari):

it's with the solution...